### v2

# Draw-In-Mind: Rebalancing Designer-Painter Roles in Unified Multimodal Models Benefits Image Editing

Abstract
In recent years, integrating multimodal understanding and generation into a single unified model has emerged as a promising paradigm. While this approach achieves strong results in text-to-image (T2I) generation, it still struggles with precise image editing. We attribute this limitation to an imbalanced division of responsibilities. The understanding module primarily functions as a translator that encodes user instructions into semantic conditions, while the generation module must simultaneously act as designer and painter, inferring the original layout, identifying the target editing region, and rendering the new content. This imbalance is counterintuitive because the understanding module is typically trained with several times more data on complex reasoning tasks than the generation module. To address this issue, we introduce Draw-In-Mind (DIM), a dataset comprising two complementary subsets: (i) DIM-T2I, containing 14M long-context image–text pairs to enhance complex instruction comprehension; and (ii) DIM-Edit, consisting of 233K chain-of-thought imaginations generated by GPT-4o, serving as explicit design blueprints for image edits. We connect a frozen Qwen2.5-VL-3B (Bai et al., 2025) with a trainable SANA1.5-1.6B (Xie et al., 2025a) via a lightweight two-layer MLP, and train it on the proposed DIM dataset, resulting in DIM-4.6B-T2I/Edit. Despite its modest parameter scale, DIM-4.6B-Edit achieves SOTA or competitive performance on the ImgEdit and GEdit-Bench benchmarks, outperforming much larger models such as UniWorld-V1 (Lin et al., 2025) and Step1X-Edit (Liu et al., 2025). These findings demonstrate that explicitly assigning the design responsibility to the understanding module provides significant benefits for image editing. Our dataset and models are available at https://github.com/showlab/DIM.
Figure 1: Upper: We employ a lightweight MLP connector to bridge a frozen MLLM, i.e., Qwen2.5-VL-3B (Bai et al., 2025), with a trainable DiT, i.e., SANA1.5-1.6B (Xie et al., 2025a), forming DIM-4.6B-Edit. In the editing process, we first leverage an external designer to produce a textual blueprint in a chain-of-thought style, which is then provided to DIM-4.6B-Edit to carry out precise image editing. Lower: DIM-4.6B-Edit establishes new state-of-the-art results on the challenging ImgEdit benchmark across diverse designers, while requiring $5\times$ fewer parameters than existing frontier models. These results highlight both the effectiveness of the proposed DIM dataset and the generalizability of our approach.
1 Introduction
Over the past few years, considerable effort has been devoted to developing unified models capable of both multimodal understanding and generation. Many such trials, e.g., Show-o (Xie et al., 2024) and MetaQuery (Pan et al., 2025), have achieved impressive results on T2I generation, yet this paradigm falters when extended to instruction-guided image editing. Even recent methods such as BAGEL (Deng et al., 2025), UniWorld-V1 (Lin et al., 2025), and Step1X-Edit (Liu et al., 2025) struggle, as evidenced by the substantial performance gap with proprietary models like GPT-4o-Image (OpenAI, 2025) on the ImgEdit and GEdit-Bench benchmarks. While much concurrent research focuses on scaling parameters and data or on architectural modifications, in this paper we identify a novel challenge underlying current image editing models: a fundamental imbalance division of responsibilities between the understanding and generation modules.
Specifically, we observe that current image editing models often translate user instructions into semantic conditions through a semantic encoder, typically a multimodal large language model, yet this process lacks intermediate reasoning or refinement. The resulting conditions are then forwarded to the generation module, which is responsible for completing the editing process. At this stage, the generation module must simultaneously infer the original layout, determine the editing region, and render the new content. In this paradigm, the understanding module functions merely as a translator, while the generation module is burdened with the demanding tasks of both design and painting. This arrangement contrasts with natural human workflows, where planning and refinement typically precede the act of drawing. A more intuitive strategy is therefore to assign design-oriented reasoning to the understanding module while allowing the generation module to focus exclusively on painting.
Motivated by this observation, we introduce Draw-In-Mind (DIM), a dataset consisting of two complementary subsets: (i) DIM-T2I that contains 14M long-context image–text pairs annotated across 21 dimensions by in-house models to lay the groundwork for complex chain-of-thought comprehension; and (ii) DIM-Edit that comprises 233K high-quality chain-of-thought imagination generated by GPT-4o from existing image editing data, enabling the model to interpret explicit design plans from an external designer. We then establish a simple baseline by concatenating a frozen MLLM, i.e., Qwen2.5-VL-3B, with a trainable DiT, i.e., SANA1.5-1.6B, via a two-layer MLP and train it end-to-end on both public data and the proposed DIM dataset, resulting in DIM-4.6B-T2I/Edit. During edit inference, we employ an arbitrary external designer, feeding its chain-of-thought imagination into the model to guide precise image edits. The framework and performance overview are illustrated in Figure 1. Despite its simplicity, DIM-4.6B-Edit matches or outperforms $5\times$ larger models such as Step1X-Edit (Liu et al., 2025) and UniWorld-V1 (Lin et al., 2025) on the ImgEdit benchmark. These results validate the effectiveness of the proposed DIM dataset and confirm the benefit of shifting the design responsibility from the generation module to the understanding module.
To summarize, we make the following contributions in this paper:
  • We pinpoint a fundamental imbalanced division of responsibilities in current image editing models, which overburdens the generation module with both design and painting tasks.
  • We introduce Draw-In-Mind (DIM), a unified dataset with two complementary subsets: DIM-T2I and DIM-Edit. This dataset explicitly frees the generation module from design responsibility and enables it to concentrate on painting, leading to substantial improvements in editing performance.
  • We establish a simple baseline by connecting a frozen Qwen2.5-VL-3B with a trainable SANA1.5-1.6B via a two-layer MLP and train it on DIM. Despite its modest size and simple architecture, DIM-4.6B-Edit outperforms $5\times$ larger competitors, validating the efficacy of DIM.
2 Related Work
2.1 Existing Image Generation Datasets
T2I Datasets. Existing T2I datasets have provided many high-quality image–text pairs. They can be roughly grouped into three categories: (i) purely AI-generated data, e.g., JourneyDB (Sun et al., 2023) and MidJourney-V6 (CortexLM, 2025), which collect images from the MidJourney API, and HQ-Edit (Hui et al., 2024), which generates images via DALL-E 3; (ii) real-world data, e.g., COCO (Lin et al., 2014), which harvests images from Flickr and annotates them by human workers; and (iii) mixed data, e.g., InstructP2P (Brooks et al., 2023), which sources real images from LAION-Aesthetics and produces edited variants via Prompt2Prompt (Hertz et al., 2022). Although these datasets deliver high perceptual quality, their prompts are typically short, limiting their utility for complex chain-of-thought reasoning in the editing stage. To ensure broad concept coverage, we opt for harvesting real-world images and annotate them with our in-house models from 21 dimensions, yielding 14M long-context image–text pairs, namely DIM-T2I, that form a robust foundation for complex CoT-guided editing.
Image Editing Datasets. Most large-scale image editing datasets either employ AI editors for end-to-end modification, e.g., InstructP2P (Brooks et al., 2023) and HQ-Edit (Hui et al., 2024), or adopt a two-stage pipeline that first localizes the edit region via grounding models and then applies inpainting to alter the target objects, e.g., UltraEdit (Zhao et al., 2024). A few efforts enlist human experts to annotate small-scale but high-quality edit pairs, e.g., MagicBrush (Zhang et al., 2023) and SEED-Data-Edit-Part3 (Ge et al., 2024). However, their instructions are typically brief and occasionally misaligned with the corresponding image pairs. In contrast, our DIM-Edit comprises 233K deliberately designed chain-of-thought imaginations derived from these existing editing datasets. These rich and detailed CoT instructions act as explicit design blueprints, lighten the cognitive load on the generation module, and significantly improve editing performance.
2.2 Unified Models for Image Generation
T2I Models. In recent years, numerous successful attempts have been made to integrate understanding and generation modules into a unified architecture. These approaches can be broadly categorized into two technical routes: (i) Integrative approaches, e.g., Show-o (Xie et al., 2024) and Janus (Wu et al., 2025a), which typically adopt an autoregressive generation paradigm to produce both image and text tokens; and (ii) Connector-based approaches, e.g., MetaQuery (Pan et al., 2025), which use a connector to bridge an understanding module and a generation module. Since the understanding and generative capabilities are tightly coupled in the former architecture and sometimes lead to conflicts that degrade both, we adopt the connector-based design to preserve state-of-the-art cognitive ability by freezing the understanding module while enhancing generation performance.
Image Editing Models. When it comes to image editing, the challenge becomes significantly harder, as neither the latest integrative models (Lin et al., 2025) nor connector-based ones (Liu et al., 2025) achieve satisfactory performance on mainstream benchmarks such as ImgEdit and GEdit-Bench compared to proprietary models like GPT-4o-Image, even when employing large-scale understanding and generation models such as Qwen2.5-VL-7B (Bai et al., 2025) and FLUX.1-dev (Labs, 2024). This suggests that simply scaling model size is not an effective strategy for improving image editing capability. In this work, we take a different approach by addressing the problem from a perspective of imbalanced division of responsibilities. We propose DIM-4.6B-Edit, which leverages an external designer to create blueprints in a CoT manner in the textual space before editing. Despite having only 1.6B generative parameters, our model achieves SOTA editing performance, highlighting the effectiveness of shifting the design responsibility to the understanding module.
3 Methodology
3.1 The Draw-In-Mind (DIM) Dataset
3.1.1 DIM-T2I
Current image generation datasets perform well in terms of prompt–image alignment and image perceptual quality. However, the prompts in existing datasets are typically short and simple, as shown in Table 1. While these prompts accurately capture the semantics of the target image, they fall short in fostering long-context comprehension, which is an essential foundation for complex CoT–guided image editing. To bridge this gap, we collect 14M images with resolutions higher than $512\times 512$ from the web and annotate them using internal models across 21 dimensions listed in Appendix B, resulting in DIM-T2I. Our produced annotations are long and detailed, thoroughly covering all dimensions. As shown in Table 1, the average prompt length is at least twice that of existing corpora, effectively establishing a strong basis for complex CoT-guided image editing.
Table 1: The statistics of existing high-quality datasets and our proposed DIM dataset. APL is short for Average Prompt Length, counted by word numbers.
Dataset Name Size Source APL
Text-to-Image
MidJourney-V6 (CortexLM, 2025) 1.2M AI Gen. 9.59
COCO (Lin et al., 2014) 0.4M Real 10.46
InstructP2P (Brooks et al., 2023) 0.6M Real & AI Gen. 11.37
JourneyDB (Sun et al., 2023) 4.2M AI Gen. 29.27
HQ-Edit (Hui et al., 2024) 0.2M AI Gen. 38.08
Dimba (Fei et al., 2024) 0.3M Real 78.29
DIM-T2I 14M Real 146.76
Image Editing
MagicBrush (Zhang et al., 2023) 8K Real 6.50
SEED-Data-Edit-Part3 (Ge et al., 2024) 82K Real 7.39
UltraEdit (Zhao et al., 2024) 4M AI Gen. 8.32
ShareGPT-4o-Image (Chen et al., 2025b) 46K AI Gen. 34.75
DIM-Edit 233K Real & AI Gen. 252.64
It is worth noting that different from existing datasets that typically apply aesthetic filtering (Fei et al., 2024; Brooks et al., 2023) or ensure perceptual quality by generating images directly with AI (Sun et al., 2023; CortexLM, 2025), our DIM-T2I is collected directly from the real world without any post-filtering, ensuring broad coverage of world knowledge. DIM-T2I demonstrates that even without complex perceptual-level filtering on the visual side, lower-quality images paired with deliberately curated long-context prompts can still serve as an effective corpus for T2I generation, evidenced by the SOTA performance of DIM-4.6B-T2I trained on this dataset.
3.1.2 DIM-Edit
As for image editing, the short-prompt issue is even more pronounced in current datasets. As shown in Table 1, prompts in mainstream datasets are generally overly simplistic, often consisting of only a few descriptive words. Such data is not conducive to effective image editing learning, as the prompts may fail to accurately reflect the actual changes between the source and target images. This phenomenon can be attributed to two main reasons: (i) Inaccurate AI editing or human misoperation. We observe that even SOTA proprietary models like GPT-4o-Image frequently over-edit images, e.g., removing objects not mentioned in the prompts. Such cases exist widely in AI-generated datasets like ShareGPT-4o-Image and UltraEdit. While in human-controlled datasets, operators may misunderstand or misapply the edits, resulting in unaligned data. (ii) Ambiguous semantics. Even if the prompt correctly describes the intended change, overly simple prompts can still result in multiple equally valid edits. For example, in SEED-Data-Edit-Part3, a common prompt is “change the background”, yet the definition of “background” varies across images, while in practice the change almost always occurs in the sky, thereby reducing the effectiveness of the resulting edit data.
Figure 2: The creation pipeline of DIM-Edit begins with a quality assessment of existing image editing data, followed by prompt optimization using GPT-4o. Finally, the optimized prompts together with the corresponding image pairs are fed into GPT-4o, which generates a four-step chain-of-thought imagination in the textual space.
In addition, existing models typically use the understanding module merely as a translator, directly converting natural language instructions into semantic conditions. The generator must then rely on these conditions to simultaneously organize the layout of the edited image, recognize existing objects, localize the edit area, render new content, and preserve unchanged regions. In other words, the generator is forced to act as both designer and painter, which is a challenging and counterintuitive setup. By contrast, humans naturally prepare a mental blueprint before editing and then simply let their hands follow it to complete the changes.
Motivated by the above issues, we propose DIM-Edit, which first optimizes prompts and then imitates human thinking to complete the edits. The DIM-Edit creation pipeline is illustrated in Figure 2. We construct it from 233K high-quality image pairs collected from three sources: (i) 160K highly consistent edit pairs from UltraEdit, referred to as UltraEdit-160K-CoT, selected using a joint SSIM, DINOv2 similarity, and CLIP similarity-based filtering; (ii) 46K semantically rich samples from the editing subset of ShareGPT-4o-Image, referred to as ShareGPT-4o-Image-CoT; and (iii) 8K human-edited images from the MagicBrush training set and 19K human-edited images from SEED-Data-Edit-Part3, specifically targeting remove operations, referred to as HumanEdit-CoT. A detailed data collection pipeline can be found in Appendix C.
After collecting raw data, we first sent the raw edit pairs to GPT-4o for prompt quality evaluation, as shown in Figure 2. The results are categorized into three groups: (i) Misaligned. The prompt does not reflect the actual edit at all, possibly due to misannotation or misoperation. (ii) Partially aligned. The target image exhibits over-editing, i.e., redundant objects are added to or removed from the source image. (iii) Aligned. The prompt fully corresponds to the edits.
Next, we take different actions to optimize the prompts based on the judgment: (i) For misaligned prompts, they are discarded outright. (ii) For partially aligned prompts, we ask GPT-4o to add details about unmentioned changes, e.g., including objects that were incorrectly removed in the prompt. (iii) For aligned prompts, we instruct GPT-4o to remove ambiguity and refine the prompt, for example, by specifying the exact objects to be edited to avoid confusion with visually similar objects.
Finally, we provide the optimized prompts, along with the source image to GPT-4o and instruct it to produce a four-step CoT imagination that emulates human editing behavior. For the sake of accuracy, we also provide the target image to it for reference. The target of each CoT step is as follows: (i) Global Layout Perception: identify and describe all key objects and their positions in the source image. (ii) Local Object Perception: describe the appearance of each object or background element in the source image, including shape, color, texture, and state. (iii) Edit Area Localization: specify which objects or regions will be modified, based on the refined instruction. (iv) Edited Image Imagination: describe the expected appearance of the edited image, with an emphasis on the modified areas. As shown in Table 1 and Figure 2, the resulting CoT imagination is not only ultra-long but also highly accurate, effectively removing the design responsibility from the generation module and thereby significantly enhancing the efficiency of image editing learning. A quality assessment of the CoTs can be found in Appendix D.
3.2 DIM-4.6B-T2I/Edit
Leveraging MLLMs to provide multimodal conditions for image generation has become a common practice recently. In this work, we first build a base T2I model and then adapt it to the editing task.
For the base T2I model, we start by establishing a simple baseline, similar to MetaQuery (Pan et al., 2025), to preserve state-of-the-art understanding capability. We select Qwen2.5-VL-3B (Bai et al., 2025) as the MLLM and SANA1.5-1.6B (Xie et al., 2025a) as the diffusion decoder for their modest size. Unlike MetaQuery, which employs a large 24-layer transformer with 1.6B parameters as a connector between the MLLM and the diffusion decoder, we adopt a much simpler design, i.e., a two-layer MLP, to directly project multimodal tokens into the generation space. We refer to this model as DIM-4.6B-T2I, illustrated in Figure 1. We train DIM-4.6B-T2I on a mixture of the proposed DIM-T2I dataset and an additional 6.9M image–text pairs from MidJourney-V6 (CortexLM, 2025), COCO (Lin et al., 2014), InstructP2P (Brooks et al., 2023), JourneyDB (Sun et al., 2023), HQ-Edit (Hui et al., 2024), and Dimba (Fei et al., 2024). During training, Qwen2.5-VL-3B remains frozen, and we finetune only the parameters of the connector and SANA1.5-1.6B. Notably, we exclude distillation datasets such as BLIP3-o-60K (Chen et al., 2025a) to avoid any risk of data leakage (Wu et al., 2025b) in the following evaluation. We adopt vanilla flow matching as the sole objective, avoiding parameter-tuning tricks to highlight data effectiveness and maintain simplicity.
Thanks to the rich world knowledge and high-quality long-context prompts in the proposed DIM-T2I dataset, the trained DIM-4.6B-T2I model provides a strong foundation for complex CoT comprehension. We then adopt a two-stage training strategy to adapt it for the editing task. In the first stage, we train DIM-4.6B-T2I on the UltraEdit (Zhao et al., 2024) dataset to develop basic editing capability. Following InstructP2P (Brooks et al., 2023), we concatenate the source image with noise along the channel dimension, as illustrated in Figure 1. In the second stage, we fine-tune the stage-one model exclusively on the proposed DIM-Edit dataset, resulting in DIM-4.6B-Edit. During inference, we employ an external designer, e.g., GPT-4o, to prepare a blueprint in the same format as DIM-Edit, except without access to the target image, ensuring alignment with real usage scenarios.
4 Experiments
4.1 Experimental Setup
During training, we use AdamW as the optimizer and keep most hyperparameters unchanged for simplicity. For DIM-4.6B-T2I, we first warm up by training only the connector for one epoch with a learning rate of $2\times 10^{-5}$, then jointly train the connector and SANA1.5-1.6B for eight epochs with the same rate and a batch size of 256. For DIM-4.6B-Edit, we set the batch size to 32, training on UltraEdit for 10 epochs at a $1\times 10^{-4}$ learning rate, then finetuning on DIM-Edit for 50 epochs at a $1\times 10^{-5}$ learning rate. During inference, GPT-4o serves as the designer unless otherwise specified.
Although the primary focus of this paper is image editing, we also evaluate DIM-4.6B-T2I on T2I generation to verify the effectiveness of the proposed DIM-T2I. Specifically, we assess semantic alignment using GenEval (Ghosh et al., 2023) scores and perceptual quality using the FID on the MJHQ-30K (Li et al., 2024). Following MetaQuery (Pan et al., 2025) and Emu3 (Pan et al., 2025), we leverage an LLM to rewrite the original prompt for GenEval evaluation. For image editing, we report scores on the recently proposed ImgEdit (Lin et al., 2025) and GEdit-Bench-EN (Liu et al., 2025) benchmarks, using GPT-4.1 for evaluation to ensure fair comparison with existing results.
4.2 Main Results
4.2.1 Text-to-Image Generation
Table 2: The text-to-image generation performance on GenEval and MJHQ-30K. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively; $\dagger$ denotes using an LLM rewriter; [Uncaptioned image] and [Uncaptioned image] denote frozen and trainable parameters, respectively.
[…]
Single Obj. Two Obj. Counting Colors Position Color Attr. Overall FID
Gen. Only
PixArt-$\alpha$ (Chen et al., 2023) 0.6B[Uncaptioned image] 0.98 0.50 0.44 0.80 0.08 0.07 0.48 6.14
SDXL (Podell et al., 2023) 2.6B[Uncaptioned image] 0.98 0.74 0.39 0.85 0.15 0.23 0.55 8.76
[…]
SD3-Medium (Esser et al., 2024) 2.0B[Uncaptioned image] 0.99 0.94 0.72 0.89 0.33 0.60 0.74 11.92
Unified
Janus (Wu et al., 2025a) 1.3B[Uncaptioned image] 0.97 0.68 0.30 0.84 0.46 0.42 0.61 10.10
[…]
Show-o (Xie et al., 2024) 1.3B[Uncaptioned image] 0.98 0.80 0.66 0.84 0.31 0.50 0.68 15.18
Show-o2-7B (Xie et al., 2025b) 7.0B[Uncaptioned image] 1.00 0.87 0.58 0.92 0.52 0.62 0.76 -
Janus-Pro-7B (Chen et al., 2025c) 7.0B[Uncaptioned image] 0.99 0.89 0.59 0.90 0.79 0.66 0.80 13.48
BAGEL (Deng et al., 2025) 14.0B[Uncaptioned image] 0.99 0.94 0.81 0.88 0.64 0.63 0.82 -
[…]
We first report T2I performance on GenEval and MJHQ-30K in Table 2. Our DIM-4.6B-T2I adopts a simple architecture with very few trainable parameters yet achieves SOTA or competitive performance, demonstrating the high data quality of DIM-T2I. For semantic alignment, DIM-4.6B-T2I shows only a small gap compared to much larger models like BAGEL (Deng et al., 2025) on GenEval. Compared with MetaQuery (Pan et al., 2025), which employs a large 1.6B-parameter connector for query learning, our model achieves nearly the same performance using only a two-layer MLP connector and naive multimodal tokens. In addition, it attains optimal perceptual quality, as evidenced by the lowest FID on the aesthetics-oriented MJHQ-30K benchmark. These results indicate that even without complex aesthetic filtering, carefully crafted long-context prompts enable robust text-to-image generation, offering a practical approach for rapid large-scale dataset creation by directly harvesting images from the web.
4.2.2 Image Editing
Table 3: The image editing performance on ImgEdit. We use GPT-4.1 for evaluation to ensure consistency with the existing results reported in UniWorld-V1. * indicates results evaluated by us using the official weights; [Uncaptioned image] and [Uncaptioned image] denote frozen and trainable parameters, respectively.
Model Params Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
MagicBrush (Zhang et al., 2023) 0.9B[Uncaptioned image] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.83
Instruct-P2P (Brooks et al., 2023) 0.9B[Uncaptioned image] 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
AnyEdit (Yu et al., 2025) 1.3B[Uncaptioned image] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
UltraEdit (Zhao et al., 2024) 2.0B[Uncaptioned image] 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
Step1X-Edit (Liu et al., 2025) 7.0B[Uncaptioned image]||12.5B[Uncaptioned image] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
BAGEL (Deng et al., 2025) 14.0B[Uncaptioned image] 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
UniWorld-V1 (Lin et al., 2025) 7.0B[Uncaptioned image]||12.0B[Uncaptioned image] 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26
Janus-4o* (Chen et al., 2025b) 7.0B[Uncaptioned image] 3.35 3.35 2.25 3.01 2.18 3.32 4.71 2.49 4.04 3.19
GPT-4o-Image (OpenAI, 2025) - 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.89 4.20
DIM-4.6B-Edit 3.0B[Uncaptioned image]||1.6B[Uncaptioned image] 4.09 3.47 2.30 4.00 3.43 3.87 4.92 2.85 4.08 3.67
The image editing performance on ImgEdit is reported in Table 3. Our DIM-4.6B-Edit shows a significant improvement over previously available open source models. In comparison with other connector-based architectures such as Step1X-Edit and UniWorld-V1, which rely on a 12B FLUX backend for generation together with a 7B multimodal large language model for condition translation, DIM-4.6B-Edit achieves superior results while maintaining both a much smaller total parameter count and a very limited number of trainable parameters.
Since DIM-Edit includes high-quality images from ShareGPT-4o-Image (Chen et al., 2025b), we also evaluate Janus-4o, which is trained on the same dataset, for reference. Janus-4o achieves only suboptimal results, indicating that the improvement comes from DIM-Edit itself, whose natural and precise edit blueprints substantially enhance editing performance. These encouraging results validate our assumption that imbalanced division of responsibilities degrades image editing, confirm the soundness of our data creation pipeline, and highlight the effectiveness of the Draw-In-Mind paradigm: assigning the design responsibility to the understanding module while allowing the generation module to focus on actual editing exclusively.
Figure 3: Green and Blue: the edits of Janus-4o and Step1X-Edit; Red: the edits of our models trained on different data corpora. All variants are tuned from the base checkpoint in Table 6.
We further demonstrate the capability of DIM-4.6B-Edit through intuitive visual comparisons of editing results on four AI-generated out-of-domain images in Figure 3. Janus-4o exhibits severe distortions despite being trained on GPT-4o-generated edit pairs, while Step1X-Edit produces less natural edits (rows 2–4) and fails in complex scenarios such as row 1, which involves manipulating multiple objects. In contrast, DIM-4.6B-Edit successfully follows the instructions to produce natural and consistent edited images. Please refer to Appendix A for more visualizations.
Table 4: The overall task-wise performance on GEdit-Bench-EN Full set. * indicates results evaluated by us. Task abbreviations: Background Change (BC), Color Alter (CA), Material Alter (MA), Motion Change (MC), PS Human (PH), Style Change (SC), Subject-Add (SA), Subject-Remove (SRM), Subject-Replace (SRP), Text Change (TC), Tone Transfer ($T$), and Average (AVG).
Model BC CA MA MC PH SC SA SRM SRP TC $T$ AVG AVG w/o TC
UniWorld-V1 (Lin et al., 2025) 4.92 6.37 4.79 1.85 4.03 5.64 7.23 6.17 5.70 1.15 5.54 4.85 5.22
Janus-4o* (Chen et al., 2025b) 4.31 5.02 4.41 2.71 4.09 5.80 4.07 1.69 3.69 2.35 3.96 3.83 3.97
Step1X-Edit (Liu et al., 2025) 7.03 6.26 6.46 3.66 5.23 7.24 7.17 6.42 7.39 7.40 6.62 6.44 6.35
DIM-4.6B-Edit 7.02 6.81 6.00 4.67 5.88 7.16 7.48 6.67 6.76 2.99 6.55 6.18 6.50
We also include overall task-wise performance on GEdit-Bench-EN in Table 4. The results reveal a similar pattern as reported in UniWorld-V1 (Lin et al., 2025): Step1X-Edit achieves notable gains in the Text Change task, whereas other models, including ours, perform less effectively due to the absence of such data in DIM-Edit. Excluding the Text Change task, DIM-4.6B-Edit beats Step1X-Edit while maintaining a compact size, underscoring the high efficacy of our CoT data. Please refer to Appendix A for full GEdit-Bench-EN results.
4.3 Ablation Study
Table 5: The ImgEdit performance w.r.t. different external designers.
Designer Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
- 3.53 3.23 2.01 3.49 1.47 3.42 4.79 2.35 3.64 3.10
Qwen2.5-VL-3B (Bai et al., 2025) 3.80 3.24 2.03 3.89 3.21 3.52 4.92 2.71 4.05 3.49
Qwen2.5-VL-7B (Bai et al., 2025) 3.95 3.35 2.25 3.85 3.31 3.57 4.88 2.81 4.02 3.55
MiMo-VL-7B (Xiaomi, 2025) 3.95 3.32 2.20 3.75 2.46 3.82 4.88 2.52 3.93 3.43
InternVL3.5-8B (Wang et al., 2025) 3.98 3.40 2.05 4.14 3.30 3.84 4.94 2.77 3.89 3.59
GLM-4.1V-9B (Hong et al., 2025) 3.95 3.27 2.23 3.90 2.64 3.81 4.92 2.23 4.02 3.44
GPT-4o (Hurst et al., 2024) 4.09 3.47 2.30 4.00 3.43 3.87 4.92 2.85 4.08 3.67
Generalizability. Although our proposed DIM-Edit is annotated with GPT-4o to ensure data quality, we show that the resulting DIM-4.6B-Edit is compatible with various external designers, as reported in Table 5. In the first row, we remove the designer and directly use the raw prompt from ImgEdit. Even under this setting, DIM-4.6B-Edit achieves performance comparable to frontier models such as BAGEL, demonstrating that high-quality CoT annotations help strengthen basic editing by mitigating prompt–edit misalignment. We then replace GPT-4o with five mainstream MLLMs as external designers, i.e., Qwen2.5-VL-3B/7B (Bai et al., 2025), MiMo-VL-7B (Xiaomi, 2025), InternVL3.5-8B (Bai et al., 2025), and GLM-4.1V-9B (Hong et al., 2025). All of them deliver strong results compared to previous state-of-the-art models in Table 3, highlighting the robustness of DIM-4.6B-Edit and the generalizability of our DIM framework. Furthermore, models equipped with external designers significantly outperform the raw-prompt setting, confirming that CoT imagination effectively reduces the burden on the generation modules and enhances overall editing quality.
Table 6: Impact of data compositions during the two training stages of DIM-4.6B-Edit on ImgEdit. Stage 2 models are tuned from checkpoint .
Data Composition Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
Stage1 Non-CoT Data
ShareGPT-4o-Image 3.35 2.74 1.93 3.05 1.95 3.16 4.91 2.00 3.70 2.98
UltraEdit-4M 3.41 3.03 1.91 2.94 1.07 3.09 3.77 2.64 2.97 2.76
+ ShareGPT-4o-Image 3.85 3.09 1.84 3.71 2.26 3.51 4.88 2.17 3.67 3.22
Stage2 CoT Data
+ ShareGPT-4o-Image-CoT 4.01 3.19 2.19 3.74 2.53 3.57 4.93 2.25 3.66 3.34
+ UltraEdit-160K-CoT 3.69 3.21 1.90 2.50 1.22 3.20 3.53 2.71 3.14 2.79
   + HumanEdit-CoT 3.63 2.99 2.01 3.01 2.64 3.11 3.73 3.03 3.01 3.02
+ DIM-Edit 4.09 3.47 2.30 4.00 3.43 3.87 4.92 2.85 4.08 3.67
Data Composition. In Table 6, we present a rigorous data composition analysis for the editing task to identify the sources of performance improvements. In the first stage, we observe that training solely on ShareGPT-4o-Image already yields a satisfactory ImgEdit score, indicating strong semantic alignment, which is consistent with the behavior of Janus-4o. However, models trained exclusively on GPT-4o-generated data tend to alter the overall layout noticeably, which is undesirable. In contrast, training on UltraEdit produces slightly lower scores but preserves better consistency between the source and target images. When combining the two datasets, performance improves significantly, as the model benefits from the semantic richness while retaining the edit consistency.
In the second stage, we finetune the checkpoint trained solely on UltraEdit. The effectiveness of our CoT data is demonstrated by comparing row 4 with row 3 in Table 6, where using the CoT version of ShareGPT-4o-Image yields a significant improvement in overall scores compared with its non-CoT counterpart. We also observe that using UltraEdit-160K-CoT alone provides only marginal gains, while the HumanEdit-CoT portion has a more notable impact due to its high edit quality, though still less pronounced than the semantically rich ShareGPT-4o-Image-CoT. When combining all three CoT components, i.e., the proposed DIM-Edit, performance improves substantially once again, indicating that UltraEdit-160K-CoT and HumanEdit-CoT are crucial for maintaining edit consistency, which is consistent with the pattern of row 3.
The visualization of three variants finetuned from the base checkpoint in Table 6 is shown in Figure 3 for intuitive analysis. The variant tuned on ShareGPT-4o-Image significantly alters the layout despite following the edit prompt, while its counterpart tuned on ShareGPT-4o-Image-CoT preserves more details, indicating that CoT imagination helps maintain editing consistency. However, using ShareGPT-4o-Image-CoT alone still produces unstable edits. In contrast, the model tuned on the full DIM-Edit dataset, i.e., DIM-4.6B-Edit, achieves the best results in both semantic alignment and edit consistency, demonstrating the effectiveness of all three data components in DIM-Edit.
Table 7: Impact of CoT compositions on ImgEdit. GLP refers to Global Layout Perception, LOP to Local Object Perception, EAL to Edit Area Localization, and EII to Edited Image Imagination.
CoT Composition Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
DIM-Edit 4.09 3.47 2.30 4.00 3.43 3.87 4.92 2.85 4.08 3.67
w/o GLP 3.85 3.29 2.06 3.91 3.24 3.55 4.80 2.79 3.92 3.49
w/o LOP 3.80 3.15 1.92 3.83 3.07 3.60 4.79 2.44 3.92 3.39
w/o EAL 3.79 3.25 1.96 3.73 2.96 3.65 4.81 2.82 3.82 3.42
w/o EII 3.77 3.22 1.82 3.88 2.96 3.61 4.78 2.55 3.58 3.35
CoT Composition. We also analyze the effect of each CoT component by individually removing it, as shown in Table 6. All components contribute positively to the performance, though their importance varies. The GLP has only a minor impact, likely because it is an easy task for the generator. In contrast, the other three CoT components, i.e., LOP, EAL, and EII, have a significant effect. LOP and EAL require the model to focus on specific regions, while EII demands complex reasoning; none of these are trivial for the generator. These findings further validate the Draw-In-Mind paradigm, which reduces the cognitive burden on the generator and thereby improves performance.
5 Conclusion
In this paper, we identify a crucial issue in existing image editing models, i.e., imbalanced division of responsibilities, where the generator is burdened with complex reasoning, leading to reduced performance. To address this, we propose the Draw-In-Mind (DIM) dataset, consisting of two parts: (i) DIM-T2I, 14M web-crawled image–text pairs with carefully crafted long-context prompts that provide a foundation for complex CoT comprehension in editing; and (ii) DIM-Edit, 233K high-quality image editing pairs with detailed and precise CoT imagination. By training on the DIM dataset and incorporating an external designer during editing, we present DIM-4.6B-Edit, which achieves SOTA or competitive performance on ImgEdit and GEdit-Bench-EN while maintaining a tiny overall and trainable parameter size. These results validate our motivation to shift the design responsibility from the generation module to the understanding module, as well as the high efficiency of our proposed CoT-guided DIM dataset.
Appendix A Additional Experiments
Table 8: The overall image editing performance on GEdit-Bench-EN. We use GPT-4.1 for evaluation to ensure consistency with the existing results reported in Step1X-Edit. * indicates results evaluated by us. SC and PQ denote Semantic Consistency and Perceptual Quality, respectively.
Model Intersection subset Full set
SC PQ Overall SC PQ Overall
Proprietary Models
Gemini (Comanici et al., 2025) 6.82 7.41 6.48 6.87 7.44 6.51
GPT-4o (OpenAI, 2025) 7.40 7.90 7.14 7.22 7.89 6.98
Doubao (Gong et al., 2025) 7.87 8.10 7.59 7.74 8.13 7.49
Open-Source Models
Instruct-P2P (Brooks et al., 2023) 3.34 6.21 3.23 3.30 6.19 3.22
MagicBrush (Zhang et al., 2023) 4.56 6.34 4.24 4.52 6.37 4.19
AnyEdit (Yu et al., 2025) 3.12 5.87 2.92 3.05 5.88 2.85
OmniGen (Xiao et al., 2025) 6.04 5.86 5.15 5.88 5.87 5.01
UniWorld-V1 (Lin et al., 2025) - - - 4.93 7.43 4.85
Janus-4o* (Chen et al., 2025b) 4.69 4.68 3.91 4.64 4.57 3.83
Step1X-Edit (Liu et al., 2025) 7.29 6.96 6.62 7.13 7.00 6.44
DIM-4.6B-Edit 6.91 6.90 6.46 6.65 6.71 6.18
Table 9: The detailed task-wise performance on GEdit-Bench-EN Full set. * indicates results evaluated by us. Task abbreviations: Background Change (BC), Color Alter (CA), Material Alter (MA), Motion Change (MC), PS Human (PH), Style Change (SC), Subject-Add (SA), Subject-Remove (SRM), Subject-Replace (SRP), Text Change (TC), Tone Transfer ($T$), and Average (AVG).
Model BC CA MA MC PH SC SA SRM SRP TC $T$ AVG AVG w/o TC
Semantic Consistency
UniWorld-V1 5.17 7.21 4.71 1.14 3.49 5.98 7.42 6.50 6.04 1.07 5.52 4.93 5.32
Janus-4o* 5.48 6.68 6.00 2.75 4.04 8.03 5.10 1.74 4.27 2.11 4.88 4.64 4.90
Step1X-Edit 8.40 7.68 7.95 3.40 5.06 8.13 7.92 6.88 8.27 7.72 7.05 7.13 7.07
DIM-4.6B-Edit 7.68 7.65 7.48 4.78 5.64 8.22 8.10 7.05 7.45 2.34 6.73 6.65 7.08
Perceptual Quality
UniWorld-V1 7.59 6.82 6.86 8.68 8.61 6.58 7.61 7.28 6.78 7.44 7.48 7.43 7.43
Janus-4o* 4.00 4.20 4.08 5.73 6.07 4.40 4.77 4.07 4.72 4.44 3.78 4.57 4.58
Step1X-Edit 6.40 6.10 5.60 7.63 8.31 6.75 7.27 7.49 6.85 7.86 6.73 7.00 6.91
DIM-4.6B-Edit 6.73 6.55 5.13 7.15 7.43 6.53 7.28 6.83 6.65 6.61 6.88 6.71 6.71
Overall
UniWorld-V1 4.92 6.37 4.79 1.85 4.03 5.64 7.23 6.17 5.70 1.15 5.54 4.85 5.22
Janus-4o* 4.31 5.02 4.41 2.71 4.09 5.80 4.07 1.69 3.69 2.35 3.96 3.83 3.97
Step1X-Edit 7.03 6.26 6.46 3.66 5.23 7.24 7.17 6.42 7.39 7.40 6.62 6.44 6.35
DIM-4.6B-Edit 7.02 6.81 6.00 4.67 5.88 7.16 7.48 6.67 6.76 2.99 6.55 6.18 6.50
Table 10: The generation configuration and inference speed of Step1X-Edit and DIM-4.6B-Edit.
Model Gen. Resolution Gen. Steps Und. Params Gen. Params VAE Rate Prompt Speed
[…]
Detailed Performance on GEdit-Bench-EN. Table 8 and 9 summarize overall and detailed task-wise performance of different models on GEdit-Bench-EN, respectively. Our DIM-4.6B-Edit ranks just behind the in-domain tester, i.e., Step1X-Edit, while surpassing all other out-of-domain competitors. Moreover, among out-of-domain testers, DIM-4.6B-Edit is the only model that consistently preserves both semantic consistency and perceptual quality. This demonstrates the effectiveness of DIM-Edit, where edits with high perceptual fidelity are precisely aligned with CoT-style imagination, thereby ensuring semantic correctness.
[…]
Figure 4: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for the add operation.
Figure 5: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for the change operation.
Figure 6: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for the remove operation.
Figure 7: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for the replace operation.
Figure 8: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for style transfer.
Visualizations of Different Editing Operations. Beyond Figure 3 in the manuscript, we further visualize the edits of Janus-4o, Step1X-Edit, and our DIM-4.6B-Edit under the operations of add, change, remove, replace, and style transfer in Figure 4567, and 8, respectively. As shown, DIM-4.6B-Edit consistently preserves the overall layout while performing natural edits. For instance, in Figure 4, Janus-4o fails to generate details of the wooden cabin, while Step1X-Edit places the chimney on the river, which is counterfactual. In contrast, our DIM-4.6B-Edit carefully adds the wooden cabin while ensuring naturalness. In Figure 5, Janus-4o fails to follow the color change instruction. Step1X-Edit changes the singer’s shirt to blue but also alters fine details such as the collar shape. By comparison, our DIM-4.6B-Edit changes the shirt to red while preserving all details, including the shadow cast by the hand. In Figure 6, both DIM-4.6B-Edit and Step1X-Edit perform successful removals, whereas Janus-4o fails to remove the seaplane. In Figure 7, only DIM-4.6B-Edit captures the semantics of “majestically” and generates a roaring lion. Finally, in Figure 8, although all three models succeed in style transfer, only DIM-4.6B-Edit captures subtle visual cues, such as the green grass in the last row, and repaints them faithfully in the edits.
Appendix B DIM-T2I Analysis Dimensions
Figure 9: The 21 analysis dimensions and corresponding prompts for DIM-T2I.
Figure 10: The 21 analysis dimensions and corresponding prompts for DIM-T2I. (Continue)
Figure 9 and 10 illustrate the 21 analysis dimensions and their corresponding prompts used in the DIM-T2I annotation process. These dimensions were carefully selected to encompass both broad world knowledge and fine-grained visual attributes, such as illumination and shadow effects, which are crucial for high-quality image generation. As a result, DIM-T2I provides not only semantically rich supervision but also consistent aesthetic guidance, forming a strong foundation for precise image editing. This is further demonstrated by the trained DIM-4.6B-T2I model, which faithfully follows user instructions at the semantic level while producing perceptually high-quality images under limited parameter budgets, achieving high GenEval scores and low FID on MJHQ-30K.
Appendix C DIM-Edit Data Collection Pipeline
As stated in Section 3.1.2, we collect raw edit data from four publicly available datasets:
  • UltraEdit (Zhao et al., 2024). To ensure editing consistency, we retain only those edit pairs that satisfy the following conditions: (i) the CLIP similarity between the source and edited images is greater than 0.9; (ii) the DINOv2 similarity is greater than 0.9; (iii) the SSIM score is greater than 0.8; and (iv) the prompt does not contain “rainbow”, since many edit pairs meeting (i)–(iii) are associated with low-quality “rainbow” edits. After filtering, we obtain roughly 160K edit pairs.
  • MagicBrush (Zhang et al., 2023). We include only 8K images from the training set to avoid potential information leakage during evaluation.
  • SEED-Data-Edit-Part3 (Ge et al., 2024). Since the “remove” operation is absent in UltraEdit, we additionally select 19K edit pairs from SEED-Data-Edit-Part3 by filtering prompts that explicitly contain “remove.”
  • ShareGPT-4o-Image (Chen et al., 2025b). We include only its 46K image-to-image subset.
By combining these collected datasets, we obtain a total of 233K raw edit pairs for the proposed DIM-Edit.
Appendix D DIM-Edit Quality Assessment
We further assess the quality of the CoTs in DIM-Edit through MLLM-powered validation. Specifically, due to API quota limitations, we randomly sample 30K edit pairs from DIM-Edit and use GPT-4.1 to evaluate the quality of the GPT-4o-annotated CoTs, categorizing them into four levels:
Figure 11: The percentage distribution of each quality level in DIM-Edit.
  • Low: The optimized edit instruction does not reflect the change between the source and edited images at all.
  • Medium: The optimized edit instruction captures the major change between the source and edited images, but the chain-of-thought contains some factual errors.
  • High: The optimized edit instruction captures the major change between the source and edited images, and the chain-of-thought contains only minor factual errors.
  • Ultra-High: The optimized edit instruction accurately captures all changes between the source and edited images, and the chain-of-thought contains no factual errors.
The percentage distribution of each quality level is shown in Figure 11. Notably, no data is categorized as “Low”, while the majority falls under the “Ultra-High” level, demonstrating the strong overall quality of DIM-Edit.